PREFACE

These Proceedings contain seven of eight lectures presented at the Spring School

“Nonlinear Analysis, Function Spaces and Applications V”

held in May 23–28, 1994, in Prague and organized by the Mathematical Institute of the Academy of Sciences of the Czech Republic in collaboration with the University of West Bohemia, Pilsen.

The School has continued the four Schools held under the same name in Horní Bradlo (1978), in Písek (1982), in Litomyšl (1986) and in Roudnice (1990) whose Proceedings were published by the TEUBNER Publishing House, Leipzig, in 1979, 1982, 1986, and 1990 as volumes 19, 49, 93, and 119, respectively, in the series “Teubner–Texte zur Mathematik”.

Analogously as in the past, the organizers concentrated on several invited survey lectures presented by a smaller number of outstanding speakers. Besides, there was also possibility for discussions and exchange of ideas, a highly attended poster session was organized as well as two seminars with short communications listed at the end of the Proceedings.

The organizers devoted this Spring School to the memory of

Prof. Svatopluk Fučík

who was the co-organizer of the first Spring School and who died 15 years ago, on May 18, 1979, at the age of 34. We would like to remember his important contribution to nonlinear analysis which was so sadly cut, and at the beginning of these Proceedings, we include a short note concerning Fučík’s work written by his pupil P. Drábek.

A list of participants can be found at the end of the volume, the total number being 95 mathematicians from the following countries:

Australia (2)  Mexico (1)
Belgium (1)  Poland (7)
Canada (3)  Portugal (1)
Czech Republic (18)  Russia (6)
Finland (2)  Slovakia (1)
Georgia (2)  Spain (6)
Germany (12)  Sweden (3)
Hungary (1)  Ukraine (3)
Italy (15)  United Kingdom (2)
Japan (3)  U.S.A. (5)

Kazakhstan (1)

The organizing committee was formed by P. Drábek, B. Koubová (Secretary), M. Krbec, A. Kufner (Chairman), B. Opic, L. Pick, and J. Rákosník. Some other mathematicians participated in the organization, from which we would like to mention P. Gurka.

Prague, June 1994 The Editors